### v1

# Federated LLMs Fine-tuned with Adaptive Importance-Aware LoRA

Abstract
Federated fine-tuning of pre-trained Large Language Models (LLMs) enables task-specific adaptation across diverse datasets while preserving data privacy. However, the large model size and heterogeneity in client resources pose significant computational and communication challenges. To address these issues, in this paper, we propose a novel Heterogeneous Adaptive Federated Low-Rank Adaptation (LoRA) fine-tuned LLM framework (HAFL). To accommodate client resource heterogeneity, we first introduce an importance-based parameter truncation scheme, which allows clients to have different LoRA ranks, and smoothed sensitivity scores are used as importance indicators. Despite its flexibility, the truncation process may cause performance degradation. To tackle this problem, we develop an importance-based parameter freezing scheme. In this approach, both the cloud server and clients maintain the same LoRA rank, while clients selectively update only the most important decomposed LoRA rank-1 matrices, keeping the rest frozen. To mitigate the information dilution caused by the zero-padding aggregation method, we propose an adaptive aggregation approach that operates at the decomposed rank-1 matrix level. Experiments on the 20 News Group classification task show that our method converges quickly with low communication size, and avoids performance degradation when distributing models to clients compared to truncation-based heterogeneous LoRA rank scheme. Additionally, our adaptive aggregation method achieves faster convergence compared to the zero-padding approach.
Index Terms:
Large Language Model, Low-Rank Adaptation, Federated Learning
publicationid: pubid: 978-1-6654-3540-6/22/\$31.00 © 2025 Crown
I Introduction
Large Language Models (LLMs) have exhibited exceptional performance in understanding and generating natural language across a wide range of tasks, including applications in chatbots and search engines[1]. To achieve optimal performance on specific tasks, these pre-trained models often require further fine-tuning. However, conventional fine-tuning methods are typically centralized, involving the collection of raw data on a single client, which raises significant privacy concerns. To address this issue, Federated Learning (FL) [2] has emerged as a promising solution by enabling collaborative model training across decentralized LLM agents, where only model weights are shared instead of raw data, thereby preserving data privacy. Integrating FL with LLMs allows for leveraging diverse data sources, resulting in more robust and generalized models.
As FL increasingly relies on wireless networks to support collaboration across clients, communication has emerged as a critical bottleneck in fine-tuning LLMs. Unlike centralized settings, FL requires frequent transmission of model updates over limited-bandwidth connections, posing significant challenges for models with billions of parameters, such as LLaMA3-8B[3]. Even in half-precision, transmitting these parameters demands approximately 16 GB per update, resulting in delays that significantly impact training efficiency. Moreover, the large parameter size imposes high demands on the client’s computational capabilities.
Parameter-Efficient Fine-Tuning (PEFT) has been proposed to alleviate the aforementioned challenges[4, 5]. One of the widely used PEFT methods is Low-Rank Adaptation (LoRA), which freezes the original pre-trained parameters of LLM and trains a smaller number of additional parameters instead[6]. This approach significantly reduces computational and storage requirements while maintaining model performance comparable to full fine-tuning. Instruction tuning, a technique to improve LLMs’ ability to follow diverse natural language instructions across various tasks, can benefit from such parameter-efficient methods. Authors in [7] proposed an approach that integrates FL with LoRA to enhance the instruction tuning of LLMs. This method enables clients to only train and optimize the LoRA layers using locally diverse instruction data, reducing privacy exposure risks by sharing only the LoRA weights, which significantly lowers the demands on communication and client computational capabilities. However, the study assumed a uniform LoRA rank across all clients and does not consider the potential impacts of client resource heterogeneity.
To accommodate heterogeneous computational capacities, the authors in [8] proposed HETLoRA that assigned different LoRA ranks to different clients. Additionally, this work introduced a zero-padding aggregation method, and a global model truncation distribution method to address the challenges posed by distinct LoRA ranks. However, the zero-padding aggregation method can dilute the information learned by high-rank clients, and the truncation approach may lead to performance loss in the models distributed to clients.
To address the challenges posed by different LoRA ranks, the authors in [9] synthesized a full-size LoRA through the direct multiplication of the A and B matrices, ensuring consistent dimensions of the full-size LoRA across clients. This full-size LoRA is then aggregated to form an updated global model, which is then decomposed using Singular Value Decomposition (SVD). The SVD components are further processed using a low-rank approximation to fit different client ranks before being distributed to the clients. However, the reconstruction method in [9] may lose information on the cross-relation across clients during aggregation, and the low-rank approximation distribution method still inevitably leads to performance loss.
In this paper, to address the limitations of previous work [7, 8, 9], we propose a novel Heterogeneous Adaptive Federated LoRA fine-tuned LLMs framework (HAFL) to handle the challenges posed by heterogeneous client resources. This framework ensures that each client’s LoRA rank is the same as the rank of the global LoRA at the cloud server for model compatibility. Each client selectively fine-tunes a portion of the LoRA layers based on its computational capabilities, keeping the remaining layers frozen rather than being truncated. Compared to HETLoRA using different ranks for different clients[8], our approach does not result in performance loss due to the truncation when adapting the highest-rank global LoRA to clients. Additionally, the freezing operation ensures that our scheme does not increase the consumption of training resources such as computational power and memory usage. Our contributions can be summarized as follows:
  • We propose a federated LoRA fine-tuned LLM framework. To accommodate client resource heterogeneity, we first introduce an importance-based parameter truncation scheme, which allows clients to have different ranks to fine-tune the more critical LoRA layers. To avoid the performance degradation resulting from the truncation process, we further develop an importance-based parameter freezing scheme. In this approach, both the cloud server and clients maintain the same high LoRA rank, while clients selectively update only the most important LoRA rank-1 matrices while keeping others frozen.
  • We propose an adaptive global model aggregation method for the cloud server. Rather than aggregating the entire LoRA matrices, we perform aggregation at the decomposed rank-1 matrix level. For each rank-1 matrix, only the clients that update and upload it will participate in the aggregation process. Unlike the zero-padding method, which allows clients with non-updated rank-1 matrices to join the aggregation process using zero values, our method effectively prevents information dilution.
  • We conduct experiments on the text classification task. The results show that our proposed HAFL method converges quickly with low communication overhead. Moreover, our importance-based partial freezing scheme effectively maintains client model accuracy during global model parameter distribution, outperforming the truncation approach. Additionally, our proposed adaptive aggregation method achieves faster convergence compared to the zero-padding approach in [8].
The rest of the paper is organized as follows: Section II introduces the system model; Section III presents our novel importance-based parameters truncation and freezing schemes and our adaptive global aggregation; Section IV presents the numerical results; and finally, Section V concludes the paper.
Figure 1: HAFL Framework
II System Model
As shown in Fig. 1, our proposed HAFL framework consists of a cloud server and a set of $N$ clients, denoted as $\mathcal{N}=\{1,2,\ldots,N\}$. Each client and the cloud server have a LoRA based LLM, where the parameters of the original pre-trained LLM are frozen, and only the incorporated low-rank decomposition matrices are trainable. Each client $k\in\mathcal{N}$ has a local dataset $\mathcal{D}_{k}=\{(x_{i},y_{i})\}_{i=1}^{D_{k}}$, where $x_{i}$ is the $i$th input data sample, $y_{i}$ is the corresponding labeled output, and $D_{k}$ is the number of data samples. Considering the limited communication resource, in each communication round, a subset of clients, $\mathcal{K}$ of size is randomly selected from the entire set of clients $\mathcal{N}$.
II-A Cloud LoRA Parameters Broadcast
The model updates of the LoRA based LLM at the cloud server and the clients can be expressed as
$\mathbf{W}=\mathbf{W}_{\text{pre}}+\Delta\mathbf{W}=\mathbf{W}_{\text{pre}}+\mathbf{B}\mathbf{A},$ (1)
where $\mathbf{W}_{\text{pre}}\in\mathbb{R}^{d\times l}$ denotes the fixed pre-trained model weights, and $\Delta\mathbf{W}\in\mathbb{R}^{d\times l}$ is the trainable matrix. The low-rank decomposition of $\Delta\mathbf{W}$ is given by $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times l}$ are the low-rank matrices with rank $r$. Here, $d$ and $l$ are the dimensions of the model’s weight matrix, and $r$ is the rank of the low-rank approximation, which is typically much smaller than both $d$ and $l$.
For different clients $k$, the LoRA rank is denoted as $r_{k}$. In our importance-based parameter truncation scheme, the client LoRA rank $r_{k}$ is selected from the range $[r_{\text{min}},r_{\text{max}}]$. In our importance-based parameter freezing scheme, the rank $r_{k}$ is set to be equal to $r_{\text{max}}$. Regardless of the scheme used, the LoRA rank of the global model at the cloud server is denoted as $r_{g}$, where $r_{g}=r_{\text{max}}$.
During each communication round, the cloud server broadcasts the LoRA parameters, which are the low-rank matrices $\mathbf{B}_{g}$ and $\mathbf{A}_{g}$, to all clients $k\in\mathcal{K}$. Taking the communication round $t$ as an example, these parameters are denoted as
$\Theta_{\text{LoRA}}^{(t)}=\{\mathbf{B}_{g}^{(t)},\mathbf{A}_{g}^{(t)}\},$ (2)
where $\mathbf{B}_{g}^{(t)}\in\mathbb{R}^{d\times r_{g}}$ and $\mathbf{A}_{g}^{(t)}\in\mathbb{R}^{r_{g}\times l}$.
II-B Local Model Training and Global Aggregation
In each communication round, the selected set of clients $\mathcal{K}$ receives the LoRA parameters $\Theta_{\text{LoRA}}^{(t)}$ broadcasted by the cloud server. Each client $k\in\mathcal{K}$ then begins local model training using its local dataset $\mathcal{D}_{k}$. In this process, clients keep the pre-trained model weights $\mathbf{W}_{\text{pre}}$ frozen. They initialize their local LoRA parameters $\mathbf{B}_{k}^{t}$ and $\mathbf{A}_{k}^{t}$ with the received $\Theta_{\text{LoRA}}^{(t)}$ and optimize these parameters by minimizing a local loss function.
After completing local training, clients upload their updated LoRA parameters $\{\mathbf{B}_{k}^{(t+1)},\mathbf{A}_{k}^{(t+1)}\}$ to the cloud server. Upon receiving updates from all clients, the cloud server aggregates these parameters to update the global LoRA parameters $\Theta_{\text{LoRA}}^{(t+1)}$, preparing for the next communication round. This approach allows the HAFL framework to effectively optimize the model using distributed data while preserving data privacy.
III Importance-based Parameters Truncation and Freezing Schemes
In this section, to handle the heterogeneity of client resources, we first introduce an importance-based parameter truncation scheme at the local LLM. To address the issue of performance degradation resulting from the truncation scheme, we subsequently introduce an importance-based parameter freezing scheme at the local LLM.
III-A Decomposed LoRA Rank-1 Matrices Importance Evaluation
In FL environments, clients often have different computational capacities, making it challenging to train all clients’ LoRA parameters at a uniform rank across all clients. To address this, the LoRA parameter product $\mathbf{B}\mathbf{A}$ will be decomposed into smaller, rank-1 matrices. As shown in Fig. 1, we take the LoRA $r_{g}=3$ as an example. The cloud server calculates an importance score for each rank-1 matrix and broadcasts this information to the clients. Clients can then prioritize these rank-1 matrices, training the most important ones first.
III-A1 Matrix Decomposition
The cloud server will decompose the product $\mathbf{B}_{g}\mathbf{A}_{g}$ into rank-1 matrices as
$\mathbf{B}_{g}\mathbf{A}_{g}=\sum_{i=1}^{r}\mathbf{b}_{i}\mathbf{a}_{i},$ (3)
where $\mathbf{b}_{i}\in\mathbb{R}^{d\times 1}$ is the $i$th column vector of $\mathbf{B}_{g}$, and $\mathbf{a}_{i}\in\mathbb{R}^{1\times l}$ is the $i$th row vector of $\mathbf{A}_{g}$.
III-A2 Element-wise Importance Calculation
To assess the significance of each element within the vectors $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$, we employ a sensitivity-based metric[10]. This involves computing the gradient of the loss function $L$ with respect to each element and using the product of the element’s value and its gradient as a measure of importance. However, in the context of FL, the global model does not have a direct loss function $L$ to calculate the gradient. Instead, we approximate the gradient by scaling the change in parameter values between consecutive communication rounds using the learning rate. Then, we define a unified importance function $I(w_{ij})$ for any trainable parameter $w_{ij}$ as follows:
$I(w_{ij})=\left|w_{ij}\cdot\frac{\Delta w_{ij}}{\eta}\right|,$ (4)
where $w_{ij}$ represents each element $b_{ji}$ in $\mathbf{b}_{i}$, with $j\in\{1,\ldots,d\}$, or each element $a_{iq}$ in $\mathbf{a}_{i}$, with $q\in\{1,\ldots,l\}$. The term $\Delta w_{ij}=w_{ij}^{(t)}-w_{ij}^{(t-1)}$ represents the change in the parameter value between two consecutive global updates, and $\eta$ is the learning rate.
However, due to the high variability and uncertainty in estimating sensitivity from mini-batch samples, it is crucial to smooth these importance scores to obtain more reliable indicators. We achieve this by applying an exponential moving average to the sensitivity scores and quantifying the uncertainty[11].
First, the smoothed sensitivity $\bar{I}^{(t)}(w_{ij})$ is calculated using an exponential moving average:
$\bar{I}^{(t)}(w_{ij})=\beta_{1}\bar{I}^{(t-1)}(w_{ij})+(1-\beta_{1})I^{(t)}(w_{ij}),$ (5)
where $\beta_{1}$ is a smoothing factor between 0 and 1.
Next, the uncertainty $\bar{U}^{(t)}(w_{ij})$ is quantified to capture the variability in the sensitivity scores:
$\bar{U}^{(t)}(w_{ij})=\beta_{2}\bar{U}^{(t-1)}(w_{ij})+(1-\beta_{2})\left|I^{(t)}(w_{ij})-\bar{I}^{(t)}(w_{ij})\right|,$ (6)
where $\beta_{2}$ is another smoothing factor between 0 and 1.
Finally, the combined importance score $s^{(t)}(w_{ij})$ is defined as the product of the smoothed sensitivity and the uncertainty:
$s^{(t)}(w_{ij})=\bar{I}^{(t)}(w_{ij})\cdot\bar{U}^{(t)}(w_{ij}).$ (7)
By incorporating these smoothing techniques, we enhance the reliability of the importance scores, making them more robust to the stochasticity in FL environments.
III-A3 Aggregation of Importance Scores
To obtain the overall importance score for the rank-1 matrix $\mathbf{b}_{i}\mathbf{a}_{i}$, we sum the smoothed and uncertainty adjusted importance scores for all elements in $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$. This approach ensures that the aggregated score reflects the reliability and significance of each parameter. The overall importance score $S_{i}$ of rank-1 matrix is calculated using:
$S_{i}=\sum_{j=1}^{d}s(b_{ji})+\sum_{q=1}^{l}s(a_{iq}),$ (8)
where $s(b_{ji})$ and $s(a_{iq})$ are the smoothed and uncertainty-adjusted importance scores for the elements of $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$, respectively. This aggregation method provides a robust measure of the importance of each rank-1 matrix in the decomposition of $\mathbf{B}\mathbf{A}$.
III-A4 Broadcast Importance Scores
After calculating the importance scores $S$ for each rank-1 matrix, we obtain a list of scores $\left[S_{1},S_{2},\ldots,S_{r}\right]$. This list represents the importance of each rank-1 matrix in the decomposition of $\mathbf{B}\mathbf{A}$. The list of importance scores is then broadcast to all clients.
III-B Importance-based Parameters Truncation Scheme
To accommodate diverse computational capabilities among clients, we allow each client to have a different LoRA rank $r_{k}\in[r_{\text{min}},r_{\text{max}}]$. However, this flexibility introduces the challenge of aligning the cloud server’s global model with the different LoRA ranks of each client’s local model. To address this issue, we propose an importance-based parameter truncation scheme.
III-B1 Truncation Process
In each communication round $t$, the cloud server broadcasts the LoRA parameters $\Theta_{\text{LoRA}}^{(t)}$ to all clients. Upon receiving these parameters, each client $k\in\mathcal{K}$ performs the following steps.
Selection of Significant rank-1 matrices
Each client $k\in\mathcal{K}$ receives the list of importance scores $[S_{1},S_{2},\ldots,S_{r}]$ from the server. Based on these scores, client $k$ selects the top $r_{k}$ rank-1 matrices as
$\mathcal{I}_{k}=\text{topk}\left([S_{1},S_{2},\ldots,S_{r}],r_{k}\right),$ (9)
where $\mathcal{I}_{k}$ represents the set of indices corresponding to the rank-1 matrices selected by client $k$.
Truncation of Low-Rank Matrices
Client $k$ then truncates the low-rank matrices $\mathbf{B}_{g}$ and $\mathbf{A}_{g}$ to align with the local model’s LoRA rank requirements. Specifically, the client retains $r_{k}$ column vectors $\mathbf{b}_{i}$ and $r_{k}$ row vectors $\mathbf{a}_{i}$ corresponding to the highest important scores:
$\mathbf{B}_{k}=\mathbf{B}_{g}[:,\mathcal{I}_{k}],\quad\mathbf{A}_{k}=\mathbf{B}_{g}[\mathcal{I}_{k},:].$ (10)
III-B2 Local Model Update
The truncated low-rank matrices $\mathbf{B}_{k}$ and $\mathbf{A}_{k}$ will be copied to the LoRA parameters of client $k$’s local model. Subsequently, client $k$ trains the local model using its local dataset $\mathcal{D}_{k}$. In the parameters truncation scheme, the local loss function for each client $k$ is defined as
$L_{k}(\mathbf{B}_{k},\mathbf{A}_{k})$ $=\frac{1}{|\mathcal{D}_{k}|}\sum_{\xi\in\mathcal{D}_{k}}\ell\left(\left(\mathbf{B}_{k},\mathbf{A}_{k}\right),\xi\mid\mathbf{W}_{\text{pre}}\right)$
$+\frac{\lambda}{2}\left(\|\mathbf{B}_{k}\|^{2}+\|\mathbf{A}_{k}\|^{2}\right),$ (11)
[…]
III-C Importance-based Parameters Freezing Scheme
The truncation process introduced above will inevitably result in performance degradation when distributing the model to clients. To address this issue, we propose a novel importance-based parameters freezing scheme, where both local models and the global model utilize the same maximum LoRA rank, denoted as $r_{k}=r_{g}=r_{\text{max}}$. To accommodate client resource constraints, clients will freeze a portion of the LoRA parameters based on their importance. We denote the freezing ratio as $\alpha_{k}$, which represents the proportion of LoRA parameters that are frozen relative to the total parameters.
III-C1 Freezing Process
In the HAFL framework, each client updates its local LoRA parameters by selectively training a subset of the parameters based on the importance scores received from the cloud server. All LoRA parameters $\Theta_{\text{LoRA}}=\{\mathbf{B}_{g},\mathbf{A}_{g}\}$ broadcasted by the server are replicated to the client’s local model, but only the most important rank-1 matrices are unfrozen and optimized by the client’s local dataset. Based on the freezing ratio $\alpha_{k}$ of each client $k$, $(1-\alpha_{k})r_{\rm max}$ rank-1 matrices will be fine-tuned.
Selection of Significant rank-1 matrices
Each client $k\in\mathcal{K}$ receives the list of importance scores $[S_{1},S_{2},\ldots,S_{r}]$ from the server. Based on these scores, client $k$ selects the top $(1-\alpha_{k})r_{\rm max}$ rank-1 matrices to actively train, where $\alpha_{k}$ can vary between clients depending on their computational capacity and resource availability. The selection is performed by choosing the indices corresponding to the highest importance scores:
$\mathcal{I}_{k}=\text{topk}\left([S_{1},S_{2},\ldots,S_{r}],(1-\alpha_{k})r_{\rm max}\right),$ (12)
where $\mathcal{I}_{k}$ is the set of indices of the selected rank-1 matrices for client $k$. The remaining rank-1 matrices’ indices are denoted as $\mathcal{U}_{k}=\{1,2,\ldots,r_{\rm max}\}\setminus\mathcal{I}_{k}$.
III-C2 Local Model Update
Each client $k$ updates its local model by minimizing a loss function $L_{k}$ over its dataset $\mathcal{D}_{k}$. The loss function is defined as
$\quad L_{k}(\mathbf{B}_{k},\mathbf{A}_{k})$
$=\frac{1}{|\mathcal{D}_{k}|}\sum_{\xi\in\mathcal{D}_{k}}\ell\left(\left(\mathbf{B}_{k}^{\mathcal{I}_{k}},\mathbf{A}_{k}^{\mathcal{I}_{k}}\right),\xi\;\middle|\;\mathbf{W}_{\text{pre}},\mathbf{B}_{k}^{\mathcal{U}_{k}},\mathbf{A}_{k}^{\mathcal{U}_{k}}\right)$
$+\frac{\lambda}{2}\left(\|\mathbf{B}_{k}^{\mathcal{I}_{k}}\|^{2}+\|\mathbf{A}_{k}^{\mathcal{I}_{k}}\|^{2}\right),$ (13)
where $|\mathcal{D}_{k}|$ represents the size of this dataset, the loss function $\ell$ evaluates the model’s performance on a data sample $\xi$. The matices $\mathbf{B}_{k}^{\mathcal{I}_{k}}=\mathbf{B}_{k}[:,\mathcal{I}_{k}]$ and $\mathbf{A}_{k}^{\mathcal{I}_{k}}=\mathbf{A}_{k}[\mathcal{I}_{k},:]$ represent the trained LoRA parameter rank-1 matrices. Additionally, $\mathbf{B}_{k}^{\mathcal{U}_{k}}=\mathbf{B}_{k}[:,\mathcal{U}_{k}]$ and $\mathbf{A}_{k}^{\mathcal{U}_{k}}=\mathbf{A}_{k}[\mathcal{U}_{k},:]$ are the untrained frozen LoRA parameter rank-1 matrices.
In the importance-based parameters freezing scheme, clients receive LoRA parameters $\{\mathbf{B}_{g},\mathbf{A}_{g}\}$ from the cloud server and replicate all of them to their local LoRA rank-1 matrices. However, only the rank-1 matrices indexed by $\mathcal{I}_{k}$ are trained and optimized, while those indexed by $\mathcal{U}_{k}$ remain frozen during local training.
III-D Adaptive Global Aggregation at the Cloud
In the HAFL framework, after completing local training, each client uploads its updated LoRA parameters to the cloud server. These parameters are then aggregated to form a new global model. For the truncation-based scheme, the upload LoRA parameters are $\mathbf{B}_{k}$ and $\mathbf{A}_{k}$. In contrast, for the importance-based parameter freezing scheme, the upload LoRA parameters are $\mathbf{B}_{k}^{\mathcal{I}_{k}}=\mathbf{B}_{k}[:,\mathcal{I}_{k}]$ and $\mathbf{A}_{k}^{\mathcal{I}_{k}}=\mathbf{A}_{k}[\mathcal{I}_{k},:]$.
Adaptive Aggregation of LoRA Parameters
The adaptive aggregation is shown in Algorithm 1, rather than aggregating LoRA parameters at the scale of the entire LoRA matrices, we perform aggregation at the decomposed rank-1 matrix level. For each rank-1 matrix, only the clients that update and upload that rank-1 matrix will participate in the aggregation process. Each client $k$ contributes to the global model based on a sparsity-based norm $z_{k}$, defined as
$z_{k}=\begin{cases}\|\mathbf{B}_{k}\mathbf{A}_{k}\|_{F},&\text{Truncation scheme}\\ \|\mathbf{B}_{k}^{\mathcal{I}_{k}}\mathbf{A}_{k}^{\mathcal{I}_{k}}\|_{F},&\text{Freezing scheme}\end{cases},$ (14)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. The total norm list $Z$ has a length of $r_{g}$, and $Z[j]$ accumulates the $z_{k}$ values from all clients that have updated the $j$th rank-1 matrix. This list is then used to normalize $z_{k}$ to obtain the contribution weight for each rank-1 matrix.
Algorithm 1 Adaptive Aggregation of LoRA Parameters
0:  LoRA parameters $\mathbf{B}_{k},\mathbf{A}_{k}$ for each client $k\in\mathcal{K}$ in parameter truncation scheme or LoRA parameters $\mathbf{B}_{k}^{\mathcal{I}_{k}},\mathbf{A}_{k}^{\mathcal{I}_{k}}$ in parameter freezing scheme.
1:  Initialize $\mathbf{B}_{g}^{(t+1)}$ and $\mathbf{A}_{g}^{(t+1)}$ to zero matrices; Initialize total norm list $Z$ with zero value
2:  for $k$ in $\mathcal{K}$ do
3:     for each $j$ in $\mathcal{I}_{k}$ do
4:        $Z[j]\mathrel{+}=z_{k}$
5:     end for
6:  end for
7:  for $k$ in $\mathcal{K}$ do
8:     Index $i=0$
9:     for each $j$ in $\mathcal{I}_{k}$ do
10:        if Truncation Scheme then
11:           $\mathbf{B}_{g}^{(t+1)}[:,j]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{B}_{k}[:,i]$
12:           $\mathbf{A}_{g}^{(t+1)}[j,:]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{A}_{k}[i,:]$
13:        else if Freezing Scheme then
14:           $\mathbf{B}_{g}^{(t+1)}[:,j]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{B}_{k}^{\mathcal{I}_{k}}[:,i]$
15:           $\mathbf{A}_{g}^{(t+1)}[j,:]\mathrel{+}=z_{k}\mathrel{/}Z[j]*\mathbf{A}_{k}^{\mathcal{I}_{k}}[i,:]$
16:        end if
17:        $i=i+1$
18:     end for
19:  end for
20:  Broadcast $\mathbf{B}_{g}^{(t+1)},\mathbf{A}_{g}^{(t+1)}$ to all clients
IV Numerical Results
In our experiments, we utilize GPT2-Large[12] as the backbone model and test on the 20 News Group classification dataset[13]. The experiment involves 100 clients and one server, with the dataset evenly distributed among these 100 clients. We assume that different clients have diverse computational capabilities, which correspond to different LoRA ranks. Specifically, in the parameter truncation scheme, 33 clients have a LoRA rank of 2, another 33 clients have a LoRA rank of 4, and the remaining 34 clients have a LoRA rank of 16. The global model at the server is configured with a LoRA rank of 16. In the parameter freezing scheme, all clients and the cloud server maintain a uniform LoRA rank of 16. However, 33 clients have a freezing ratio of 0.875, another 33 clients have a freezing ratio of 0.75, and the remaining 34 clients have a freezing ratio of 0. To simulate a real-world FL setting, 10 clients are randomly selected to participate in the training during each communication round. The hyper-parameters related to the experiment are shown in Table I.
TABLE I: Experiment Hyper-parameters
Parameter Value Parameter Value
LoRA Dropout 0.1 LoRA Bias None
LoRA Scaling Factor 32 LoRA Initialization Gaussian
LoRA Target Module attn.c_attn Optimizer Adam
Learning Rate $\eta$ 0.001 Weight Decay $\lambda$ 0.001
Token Padding Length 256 Token Truncation True
Smoothing $\beta_{1}$, $\beta_{2}$ 0.85, 0.85 Seed 0, 1, 42
Figure 2: Accuracy comparison on 20 News Group Classification
Fig. 2 presents the classification accuracy of the global model on the 20 News Group classification test set as the number of communication rounds increases. Our methods, Importance-based Parameters Freezing Adaptive Aggregation LoRA (IFALoRA) and Importance-based Parameters Truncation Adaptive Aggregation LoRA (ITALoRA), are shown with fast convergence. We also compared three other algorithms: (1) Homogeneous LoRA (HomLoRA) $r=2$, where all clients and cloud server use LoRA with a rank of 2; (2) HomLoRA $r=16$, where all clients and cloud server use LoRA with a rank of 16; and (3) IFZLoRA, which utilizes zero-padding aggregation compared to IFALoRA. The results indicate that HomLoRA $r=2$ fails to converge, and IFZLoRA converges slowly with significantly lower accuracy than our TrALoRA and IFALoRA. Our methods achieve accuracy close to that of HomLoRA $r=16$, where all clients use the highest rank.
TABLE II: Evaluation results on 20 News Group classification
Model Method Commu. Size 20NEWS(Acc.)
50 100
GPT2-Large IFZLoRA 52.28M 36.43±7.6 62.83±1.3
HomLoRA r=16 113M 67.41±0.6 69.74±0.2
HomLoRA r=2 14M 6.25±1.0 5.66±0.7
ITALoRA 52.28M 65.30±1.8 67.31±0.3
IFALoRA 52.28M 65.93±0.8 67.55±0.8
The Table II presents the performance of five different methods on the 20 News text classification task after 50 and 100 communication rounds. The values are averaged over experiments with different random seeds, and the table also records the average total parameter size uploaded by 10 clients per communication round. Specifically, HomLoRA with rank 16 achieves the highest accuracy but requires the largest communication size. In contrast, IFALoRA and ITALoRA manage to achieve nearly comparable accuracy with significantly smaller communication sizes, demonstrating a good balance between performance and efficiency. HomLoRA with rank 2 shows poor performance, indicating a lack of convergence, while IFZLoRA improves over HomLoRA r = 2 but still lags behind ITALoRA and IFALoRA in terms of accuracy.
Fig. LABEL:fig:Performance_degradation LABEL:fig:ITALoRA shows the performance degradation of the ITALoRA method when the global model is truncated to smaller client ranks. As the truncation increases, the model’s accuracy decreases, indicating a loss of information. In contrast, our IFALoRA method in Fig. LABEL:fig:Performance_degradation LABEL:fig:IFALoRA uses a freezing strategy, eliminating the need for truncation. This allows each client to maintain the same accuracy performance level as the global model. Additionally, both IFALoRA and ITALoRA have the same parameter size during training, ensuring communication efficiency while IFALoRA avoids the performance loss associated with truncation.
V Conclusion
In this paper, we proposed a HAFL framework to address privacy concerns in fine-tuning LLMs. To accommodate client resource heterogeneity, we first introduced an importance-based parameter truncation scheme at the local LLM. However, the truncation process inevitably results in performance degradation. To tackle this, we further developed an importance-based parameter frozen scheme at the local LLM. At the cloud server, we proposed an adaptive global model aggregation method to counter the information dilution problem caused by the zero-padding aggregation method. Our experimental results demonstrate that our method achieves rapid convergence with low communication overhead. Compared to the truncation approach, our importance-based partial freezing scheme maintains client model performance during global model parameter distribution. Furthermore, the adaptive aggregation method achieves faster convergence compared to the zero-padding approach.